OpenStudy (anonymous):

two fair dice are rolled repeatedly until the total number of spots that comes up is divisible by 3. what is the probability that this first happens on a roll number divisble by 3

OpenStudy (anonymous):

you probably have to sum a geometric series

OpenStudy (anonymous):

if it first happens on the third roll, probability is \[\frac{ 2 }{ 3 }*\frac{ 2 }{ 3 }*\frac{ 1 }{ 3 }=\frac{ 4 }{ 27 }\]

OpenStudy (anonymous):

So since the probability it is divisible by 3 is 1/3 and the probability it is not divisible by 3 is 2/3 so that particular sequence represents "not , not, is"

OpenStudy (anonymous):

can you explain why it is 1/3?

OpenStudy (anonymous):

Yep P(sum divisible by 3) = 1/3

OpenStudy (anonymous):

^ she's wrong XDD

OpenStudy (anonymous):

So then the sequence for getting it for the first time on the sixth roll is \[\frac{ 2 }{ 3 }*\frac{ 2 }{ 3 }*\frac{ 2 }{ 3 }*\frac{ 2 }{ 3 }*\frac{ 2 }{ 3 }*\frac{ 1 }{ 3 }\] and so on so the series will end up looking like\[\frac{ 4 }{ 27 }+\frac{ 4 }{ 27 }*\frac{ 8 }{ 27 }+\frac{ 4 }{ 27 }*(\frac{ 8 }{ 27 })\]

OpenStudy (anonymous):

shut up @sourwing XD XD

OpenStudy (anonymous):

\[(\frac{ 8 }{ 27 })^{2}\]

OpenStudy (anonymous):

sorry mistake there and then +... etc

OpenStudy (anonymous):

^ she makes alot of mistake XDD

OpenStudy (anonymous):

hence she's wrong

OpenStudy (anonymous):

you can add this up with \[\frac{ a }{ 1-r }\] with \[a=\frac{ 4 }{ 27 }\] and \[r=\frac{ 8 }{ 27 }\] and I'm gonna ring your neck if you don't shut up XD

OpenStudy (anonymous):

and she needs some beer which I will gladly give ahaha

OpenStudy (anonymous):

I got \[\frac{ 4 }{ 19 }\] nah you're making me need some brandy... or maybe a migraine pill

OpenStudy (anonymous):

@shanna13 are you getting this or would you like me to better explain it?

OpenStudy (anonymous):

you can explain better haha

OpenStudy (anonymous):

@sourwing get me a bottle of Malheur 12 and maybe I will

OpenStudy (anonymous):

but he left lol

OpenStudy (anonymous):

you probably need 2 Malheur 12 and .... a can of beer XDDD

OpenStudy (anonymous):

yeah and a shot of tequila because this is killing me @sourwing